If the 0.100-mm-diameter tungsten filament in a light bulb is to have a resistance of 0.300 \OmegaΩ at 20.0 ^o


o


C , how long should it be?

1
Expert's answer
2022-03-13T18:50:00-0400

The length is given as follows:


"l = \dfrac{AR}{\rho}"

where "A = \pi d^2/4" is the area of the filament cross-section, "d = 0.1mm = 0.1\times 10^{-3}m" is the diameter, "R = 0.3\Omega" is the resistance, and "\rho = 5.6\times 10^{-8}\Omega \cdot m" is electrical resistivity of tungsten. Thus, obtain:


"l = \dfrac{\pi d^2R}{4\rho}\\
l = \dfrac{\pi \cdot (0.1\times 10^{-3})^2\cdot 0.3}{4\cdot 5.6\times 10^{-8}} \approx 0.0421m = 4.21cm"

Answer. 4.21 cm.